a sltwater solution of 40 cups is 30%salt.Approximately how many cups of another saltwater solution of 65% salt mixture need to be added to obtain a 50% salt solution ?
Answer by ptaylor(2198)


Let x=amount of 65% solution needed
Now we know that the amount of pure salt in the 30% solution (0.30*40) plus the amount of pure salt in the 65% solution (0.65x) has to equal the amount of pure salt in the final mixture (0.50(40+x)). So our equation to solve is:
0.30*40+0.65x=0.50(40+x) get rid of parens (distributive law)
12+0.65x=20+0.50x subtract 0.50x and also 12 from each side
12-12+0.65x-0.50x=20-12+0.50x-0.50x collect like terms
0.15x=8 divide each side by 0.15
x=53 1/3 cups or approximately 53 cups-----------ans

CK
0.30*40+0.65*53=0.50*93
12+34.45=46.5
46.45~~~~~~~~~~~~~~~~~~~46.5
Hope this helps---ptaylor